OpenStudy (anonymous):

you are a gymnast of mass 72 kg, initially hanging at rest from a bar. you let go of the bar and fall vertically 92cm to the floor below. upon landing, you bend your knees, bringing yourself to rest over a distance of 35cm. the floor exerts a constant force on your body as you slow down. Determine your speed at impact, and the magnitude of the force the floor exerts on you as you slow down.

OpenStudy (anonymous):

So recall that the gravitational energy is converted to kinetic energy. so, (72kg)(9.81 m/s^2)(0.92m) = 1/2(72kg)(v^2) so solve for v = squareroot**[ 649.152/36) = 4.246 m/s down. That is the speed at impact. @genius12 could you do the rest I have to go eat. ..